# Evaluating LLMs’ Mathematical and Coding Competency through Ontology-guided Interventions

### v4

@@ Appendix B Dataset Details
In particular, there are a total of 5 maths questions for each category except Change Subject and Reverse Engineering, which have 3 and 4 questions, respectively, in More. Likewise, all but Reverse Engineering perturbation—with 4 questions—have 5 coding questions in Core.
@@ C.1 Model Details
we use version "2023-09-01-preview" for both GPT-4 and GPT-3.5.
Llama3-Instruct https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct
MetaMath-70B-V1.0 https://huggingface.co/meta-math/MetaMath-70B-V1.0
CodeLlama-70B-Instruct https://huggingface.co/codellama/CodeLlama-70b-Instruct-hf
@@ Appendix D Fine-Grained Analysis
As illustrated in Table 3, the introduction of perturbed questions poses significant challenges to all models in both maths and coding contexts. Specifically, GPT-4’s accuracy decreased from 100% to 74.2% and from 80% to 56.7% in math and coding scenarios respectively. This trend of performance degradation is even more pronounced in other LLMs, with all experiencing a decline exceeding 30 points in their weighted average performance across both the mathematical and coding datasets. For instance, GPT-3.5 witnessed a dramatic performance reduction from 80% to 35.75% on the mathematical dataset and from 80% to 47.09% for the coding dataset.
@@
(Level II) Target-wise Performance. Following Table 3, LLMs generally showed better results on logic alteration questions, which involve concrete reasoning steps in problem-solving. Despite this, even the state-of-the-art models struggled with certain perturbed versions of these questions. This indicates that while current models may possess general task-solving skills and abstract reasoning ability, there is still a limitation in their reasoning robustness when faced with altered logic. On the other hand, concept analysis questions, which demand a deeper understanding of mathematical concepts and problem-solving frameworks, resulted in lower success rates. This suggests that while current models can find correct answers, they may lack a systematic logical framework for problem-solving and struggle with analyzing and understanding different concepts necessary to answer the question.
@@ Appendix E Benchmark Difficulty Evaluation
Granularity Adjustment | Remove Constraint | 100 | 84 | -70 | ![]() |
Partial Solution | 100 | 70 | -40 | ![]() |
Solution Plan | 100 | 76 | -50 | ![]() |
Detail Expansion | 100 | 70 | -50 | ![]() |
Reasoning Adjustment | Add Restriction | 100 | 22 | +100 | ![]() |
Subsequent Question | 100 | 34 | +50 | ![]() |
Concurrent Question | 100 | 36 | +150 | ![]() |
Change Question | 100 | 42 | -70 | ![]() |
Info Recombination | 87 | 28 | +40 | ![]() |
Domain Knowledge | 80 | 56 | +450 | ![]() |
Complex Reality | 100 | 32 | +100 | ![]() |
General Solution | 100 | 24 | +0 | ![]() |
Computation Adjustment | Computation Demand | 100 | 36 | +20 | ![]() |
Change Value | 100 | 56 | -10 | ![]() |
Change Operation | 100 | 66 | +0 | ![]() |
Formulation Adjustment | Symbolic Response | 100 | 42 | +100 | ![]() |
Value Relationship | 93 | 20 | +100 | ![]() |
Variable Group | 100 | 24 | +140 | ![]() |
Backward Reasoning | 100 | 26 | +100 | ![]() |
Counterfactual | 100 | 18 | +160 | ![]() |
Solve Value | 100 | 28 | +140 | ![]() |
Identify Range | 93 | 26 | -40 | ![]() |
Question Understanding | Inherent Premise | 100 | 38 | +160 | ![]() |
Complete Missing | 100 | 60 | -50 | ![]() |
Question Formulation | 93 | 50 | +200 | ![]() |
Add Misinformation | 100 | 68 | +50 | ![]() |
Solution Understanding | Optimize Solution | 100 | 50 | +160 | ![]() |
Step Functionality | 100 | 42 | +100 | ![]() |
Theoretical Basis | 100 | 62 | -50 | ![]() |
Cost Analysis | 100 | 58 | +50 | ![]() |
Critical Thinking | Seek Clarification | 80 | 26 | -50 | ![]() |
Conditional Analysis | 93 | 16 | +200 | ![]() |
Conflicting Information | 100 | 8 | +50 | ![]() |
Surface Error | 100 | 44 | +50 | ![]() |
Hidden Error | 93 | 30 | +200 | ![]() |
Format Change | Setting Rephrase | 100 | 50 | +0 | ![]() |
Change Sequence | 100 | 52 | +0 | ![]() |
Close Format | 93 | 36 | +20 | ![]() |
Data Restructuring | 100 | 40 | +160 | ![]() |
Format Comparison | Identical Problem | 87 | 42 | +20 | ![]() |
Format Constraint | Reasoning Format | 100 | 30 | +200 | ![]() |
Reasoning Style | 100 | 34 | +170 | ![]() |
Alternative Answer | 100 | 28 | +60 | ![]() |
New Rule | 87 | 36 | +250 | ![]() |
@@
![[Uncaptioned image]](extracted/5694788/images/up-sign.png)
![[Uncaptioned image]](extracted/5694788/images/equal-sign.png)
@@ F.1 Question 1
[backgroundcolor=pink!20] A merchant wants to make a choice of purchase between 2 purchase plans: jewelry worth $5,000 or electronic gadgets worth $8,000. His financial advisor speculates that the jewelry market will go up 2.5% while the electronic gadgets market will rise 1.2% within the same month. If the merchant is looking to maximize profit at the end of this month by making a choice, how much profit would this be?
@@ F.2 Question 2
[backgroundcolor=pink!20] Question 2: John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. What is the total inner volume of all 3 boxes?
@@ F.3 Question 3
[backgroundcolor=pink!20] Question 3: Kylar went to the store to buy glasses for his new apartment. One glass costs $5, but every second glass costs only 60% of the price. Kylar wants to buy 16 glasses. How much does he need to pay for them?
@@ F.4 Question 4
[backgroundcolor=pink!20] Question 4: Vicki is planning a pop concert at her high school. The show will be 2 hours. She is allowing each group 2 minutes to get on stage, 6 minutes to perform, and then 2 minutes to exit the stage. If she allows a 10-minute intermission, how many groups can perform in the concert?
@@ F.5 Question 5
[backgroundcolor=pink!20] Question 5: Together Lily, David, and Bodhi collected 43 insects. Lily found 7 more than David. David found half of what Bodhi found. How many insects did Lily find?
@@ I.1 Principles behind Ontology
Solution Plan: Besides the original question, provide a high level plan of how the question should be answered, the solution will only need to execute the abstract plan. Solution Plan (Math): Provide the original question along with its mathematical structure (problem strategy) to the question, ask the model to solve the question by following the strategy. Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. What is the total inner volume of all 3 boxes? Follow this plan to solve the question: [#Solution Plan#] Subtract the thickness of the walls from each dimension of the box to get the inner dimensions. Determine the width, height, and depth of the inner box. Calculate the inner volume of one box by multiplying the width, height, and depth. Calculate the total inner volume by multiplying the inner volume of one box by the number of boxes. Solution Plan (Code): Provide the high level plan of how the code need to be written along with the question. Changed from G.1: Generate a python function that fulfills the requirement in docstring and examples usages below. You should follow the solution plan when solving the problem.
@@
Detail Expansion: Besides the original question, provide a few key important details or explanations without which is hard to solve the question. Detail Elaboration: Provide original question along with the toolbox (commonsense knowledge) to solve the question. Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes. Each box has outside dimensions of 5 inches by 6 inches by 4 inches. The walls of each box are 1 inch thick, uniformly throughout each face of the box, thereby reducing the inner dimensions of each box. The material of the boxes is uniformly distributed and does not bulge or cave in thereby affecting the inner volume. There are no internal structures or partitions inside the boxes that could further reduce the inner volume. What is the total inner volume of all 3 boxes? Example Detail: Besides providing the input and output of each example, it also provide a step by step explanation of how the input is transformed to the output. Changed from G.3: Generate a python function that fulfills the requirement in docstring and examples usages below.
@@
Restrict Question: Adding a new piece of information that serves as a constraint or modifier on the query. Example: Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes, each of which is externally measured as 5 inches by 6 inches by 4 inches. The boxes have walls that are 1 inch thick. There is also an added wooden board divider in the middle across the smallest dimension which is also 1 inch thick. What is the total inner volume of all the boxes? Restrict Requirement: Add a piece of information that serves as a constraint or modifier on the function. Changed from G.1
@@
Subsequent Question: Adding an additional query or requirement based on the answer of of the original question. Further Question: Adding an additional query that will need extra steps of calculation based on the final answer of the original query. {mdframed}[backgroundcolor=pink!20] Changed from F.2: John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. What is the total inner volume of all 3 boxes? If John wants to entirely fill these boxes with small cubes each measuring 0.5 inches on all sides, then how many cubes will he need? Further Requirement: Adding an additional requirement of transformation based on the output of the original function.
@@
Concurrent Question: Adding an additional query or requirement that is independent from the original question. Parallel Question: Adding an additional query along with the original query based on the information given in the question, the added query should inquiry a value that is irrelevant of the original answer. Example: Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. What is the total inner volume of all 3 boxes? What is the total volume of the material used to build the boxes? Parallel Requirement: Adding an additional requirement in Docstring that does not rely on the output of the original question. Changed from G.1:
@@
Change Question: Change the current query or requirement to a different but similar one based on the existing information provided inside the question. Change Query: Change the query to ask for another value that requires more computation based on the information given in the question. Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. What is the total outer volume of all 3 boxes? Change Docstring: Change the Docstring to another requirement based on the input given in the question header. Changed from G.3:
@@
Info Recombination: Combine the fundamental concepts or frameworks from another question with the original question. Info Recombination (Math): Graft mathematical structure from another question and combine with the original question. Changed from F.4: {mdframed}[backgroundcolor=pink!20] Vicki and James are planning an event at their high school that combines a pop singing concert and dance events. The whole event will be 2 hours long. Vicki is allowing each musical group 2 minutes to get on stage, 6 minutes to perform, and then 2 minutes to exit the stage. James will also perform two solo dance routines, each lasting five minutes. Considering a 10-minute intermission during the show, how many musical groups can perform at the concert? Info Recombination (Code): Merge the requirement from another coding question with existing question. Changed from G.1:
@@
Domain Knowledge: Introduce a specific knowledge in math or code and merge it with the question. Theoretical Challenge: Incorporate a specific theorem into the question so that perturbed question requires a new toolbox to solve. Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has an infinite number of boxes numbered as first, second, third, and so on. The first box is 5 inches by 6 inches by 7 inches in size. Starting from the second box each box is half the size of the previous box in each dimension. What is the total volume all the boxes combined? Code Import: The requirement requires to use a specific python library to solve the problem. Changed from G.1: Rewrite the function below to take in batch input parameters and use the multicore cpu for efficiency.
@@
Complex Reality: Add an aspect of complexity in the real world scenario. Value Probability:Introduce concept of uncertainty to deterministic values and calculate the estimation. The perturbed question will require toolbox (knowledge) of probability. Changed from F.1: {mdframed}[backgroundcolor=pink!20] A merchant wants to make a choice of purchase between 2 purchase plans: jewelry worth $5,000 or electronic gadgets worth $8,000. His financial advisor speculates that the jewelry market has a 70% chance to go up 2.5% and a 30% chance to remain the same, while the electronic gadgets market will rise 1.2% within the same month. If the merchant is looking to maximize profit at the end of this month by making a choice, how much estimated profit would this be? Example Boundary: Add boundary examples along with the existing examples. The boundary examples contains input that does not met requirement specified in the docstring. Changed from G.3: Write a function to fulfill the requirement and all the examples inside the docstring
@@
General Solution: Provide the solution in a more general scenario. Code Implementation: Develop a code function to solve the question in general. Changed from F.2: {mdframed}[backgroundcolor=pink!20] # Original Examples # Can you write a Python code to find out what is the total inner volume of all 3 boxes? Higher Order: Write a higher order function that can solve the coding problem in general. Changed from G.2
@@
Value Big: Significantly increasing the magnitude of values that pose a challenge for calculations. Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3000 boxes. Each box is 500 inches by 600 inches by 400 inches. The walls are 100 inches thick. What is the total inner volume of all the boxes? Generalize Parameter: Extend the current parameter into different python object types Changed from G.2:
Change Value: Change the content of the value to a different one. Change Subject: If there are multiple mentions in the question, Exchange values of names or references in the question. Changed from F.5: {mdframed}[backgroundcolor=pink!20] Together David, Bodhi, and Lily collected 43 insects. David found 7 more than Bodhi. Bodhi found half of what Lily found. How many insects did Lily find? Parameter Content: Change the format or meaning of the input parameter. Changed from G.3:
Change Operation: Change one operation regarding how the Values are processed. Change Calculation: Change no more than 3 words in original question so that the toolbox (mathematical operations) involved in the calculation are changed. Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes. The inner dimension of each box is 3 inches by 4 inches by 2 inches. The walls are 0.5 inches thick. What is the total outer volume of all 3 boxes? Variable Type: Change the python object type of the original parameter while keep its content the same, also specify the return variable to be in a specific type. Changed from G.3:
Symbolic Manipulation: This dimension test the abstract reasoning ability of under the same logical structure of the original question. This dimension focus on solving the general version of the original reasoning problem, rather than focus on to get a standard solution. For math, We change the context to include one or more symbolic variables to replace its original values.

### v5

@@ Appendix B Dataset Details
In particular, there are a total of 5 math questions for each category except Change Subject and Reverse Engineering, which have 3 and 4 questions, respectively, in GSMore. Likewise, all but Reverse Engineering perturbation—with 4 questions—have 5 coding questions in HumanEval-Core.
@@ C.1 Model Details
We use o1-preview with model version dated "2024-08-06".
We use GPT-4o with model version dated "2024-08-06".
We use GPT-4 with model version named "1106-Preview".
For GPT-3.5, we use version "GPT-3.5-turbo-0613".
We use Gemini-1.5-pro for Gemini model.
Llama3-Instruct https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct.
MetaMath-70B-V1.0 https://huggingface.co/meta-math/MetaMath-70B-V1.0.
CodeLlama-70B-Instruct https://huggingface.co/codellama/CodeLlama-70b-Instruct-hf.
@@ Appendix D Fine-Grained Analysis
As illustrated in Section 5.3, the introduction of perturbed questions poses significant challenges to all models in both math and coding contexts. Specifically, GPT-4’s accuracy decreased from 100% to 74.2% and from 80% to 56.7% in math and coding scenarios respectively. This trend of performance degradation is even more pronounced in other LLMs, with all experiencing a decline exceeding 30 points in their weighted average performance across both the mathematical and coding datasets. For instance, GPT-3.5 witnessed a dramatic performance reduction from 80% to 35.75% on the mathematical dataset and from 80% to 47.09% for the coding dataset.
@@
(Level II) Target-wise Performance. Following Section 5.3, LLMs generally showed better results on logic alteration questions, which involve concrete reasoning steps in problem-solving. Despite this, even the state-of-the-art models struggled with certain perturbed versions of these questions. This indicates that while current models may possess general task-solving skills and abstract reasoning ability, there is still a limitation in their reasoning robustness when faced with altered logic. On the other hand, concept analysis questions, which demand a deeper understanding of mathematical concepts and problem-solving frameworks, resulted in lower success rates. This suggests that while current models can find correct answers, they may lack a systematic logical framework for problem-solving and struggle with analyzing and understanding different concepts necessary to answer the question.
@@ Appendix E Benchmark Difficulty Evaluation
| Granularity Adjustment | Remove Constraint | 100 | 84 | -70 | ||
| Partial Solution | 100 | 70 | -40 | |||
| Solution Plan | 100 | 76 | -50 | |||
| Detail Expansion | 100 | 70 | -50 | |||
| Reasoning Adjustment | Add Restriction | 100 | 22 | +100 | ||
| Subsequent Question | 100 | 34 | +50 | |||
| Concurrent Question | 100 | 36 | +150 | |||
| Change Question | 100 | 42 | -70 | |||
| Info Recombination | 87 | 28 | +40 | |||
| Domain Knowledge | 80 | 56 | +450 | |||
| Complex Reality | 100 | 32 | +100 | |||
| General Solution | 100 | 24 | +0 | |||
| Computation Adjustment | Computation Demand | 100 | 36 | +20 | ||
| Change Value | 100 | 56 | -10 | |||
| Change Operation | 100 | 66 | +0 | |||
| Formulation Adjustment | Symbolic Response | 100 | 42 | +100 | ||
| Value Relationship | 93 | 20 | +100 | |||
| Variable Group | 100 | 24 | +140 | |||
| Backward Reasoning | 100 | 26 | +100 | |||
| Counterfactual | 100 | 18 | +160 | |||
| Solve Value | 100 | 28 | +140 | |||
| Identify Range | 93 | 26 | -40 | |||
| Question Understanding | Inherent Premise | 100 | 38 | +160 | ||
| Complete Missing | 100 | 60 | -50 | |||
| Question Formulation | 93 | 50 | +200 | |||
| Add Misinformation | 100 | 68 | +50 | |||
| Solution Understanding | Optimize Solution | 100 | 50 | +160 | ||
| Step Functionality | 100 | 42 | +100 | |||
| Theoretical Basis | 100 | 62 | -50 | |||
| Cost Analysis | 100 | 58 | +50 | |||
| Critical Thinking | Seek Clarification | 80 | 26 | -50 | ||
| Conditional Analysis | 93 | 16 | +200 | |||
| Conflicting Information | 100 | 8 | +50 | |||
| Surface Error | 100 | 44 | +50 | |||
| Hidden Error | 93 | 30 | +200 | |||
| Format Change | Setting Rephrase | 100 | 50 | +0 | ||
| Change Sequence | 100 | 52 | +0 | |||
| Close Format | 93 | 36 | +20 | |||
| Data Restructuring | 100 | 40 | +160 | |||
| Format Comparison | Identical Problem | 87 | 42 | +20 | ||
| Format Constraint | Reasoning Format | 100 | 30 | +200 | ||
| Reasoning Style | 100 | 34 | +170 | |||
| Alternative Answer | 100 | 28 | +60 | |||
| New Rule | 87 | 36 | +250 | |||
@@ F.1 Question 1
[backgroundcolor=pink!20]
A merchant wants to make a choice of purchase between 2 purchase plans: jewelry worth $5,000 or electronic gadgets worth $8,000. His financial advisor speculates that the jewelry market will go up 2.5% while the electronic gadgets market will rise 1.2% within the same month. If the merchant is looking to maximize profit at the end of this month by making a choice, how much profit would this be?
@@ F.2 Question 2
[backgroundcolor=pink!20]
Question 2: John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. What is the total inner volume of all 3 boxes?
@@ F.3 Question 3
[backgroundcolor=pink!20]
Question 3: Kylar went to the store to buy glasses for his new apartment. One glass costs $5, but every second glass costs only 60% of the price. Kylar wants to buy 16 glasses. How much does he need to pay for them?
@@ F.4 Question 4
[backgroundcolor=pink!20]
Question 4: Vicki is planning a pop concert at her high school. The show will be 2 hours. She is allowing each group 2 minutes to get on stage, 6 minutes to perform, and then 2 minutes to exit the stage. If she allows a 10-minute intermission, how many groups can perform in the concert?
@@ F.5 Question 5
[backgroundcolor=pink!20]
Question 5: Together Lily, David, and Bodhi collected 43 insects. Lily found 7 more than David. David found half of what Bodhi found. How many insects did Lily find?
@@ I.1 Principles behind Ontology
Docstring: Defines the requirement for the final output. For example, Return the greatest common divisor of two integers
Values: The type and structure of input arguments. In the above example, a (integer type) and b (integer type)
Examples: Demonstrations of how the function is used. In the case above,
@@
Partial Solution: The answer only need to solve parts of the original question
Median Inquiry: Change the original query to ask one of the intermediate values that is used to solve for the final answer of the original query. The aim is to simplify the question. Example:
Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. What is the inner volume of one box? Helper Function: Provide a helper function alongside the coding question that achieves partial function in Code Structure
Changed from Section G.1: Generate a python function that fulfills the requirement in docstring and examples usages below. You should complete the function using helper function.1def helper_function(char: str) -> str:2 """Checks if a given character is uppercase or lowercase, and flips its case."""34 if char.isupper():5 return char.lower()6 elif char.islower():7 return char.upper()8 else:9 return char1011def flip_case(string: str) -> str:1213 """For a given string, flip lowercase characters to uppercase and uppercase to lowercase by using the helper function above to achieve the requirement14 >>> flip_case(’Hello’)15 ’hELLO’16 """1718"""" -
Solution Plan: Besides the original question, provide a high level plan of how the question should be answered, the solution will only need to execute the abstract plan.
Solution Plan (Math): Provide the original question along with its mathematical structure (problem strategy) to the question, ask the model to solve the question by following the strategy.
@@
Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. What is the total inner volume of all 3 boxes? Follow this plan to solve the question: [#Solution Plan#] Subtract the thickness of the walls from each dimension of the box to get the inner dimensions. Determine the width, height, and depth of the inner box. Calculate the inner volume of one box by multiplying the width, height, and depth. Calculate the total inner volume by multiplying the inner volume of one box by the number of boxes. Solution Plan (Code): Provide the high level plan of how the code need to be written along with the question. Changed from G.1: Generate a python function that fulfills the requirement in docstring and examples usages below. You should follow the solution plan when solving the problem.1def flip_case(string: str) -> str:2 """3 Inverts the case of each character in the provided string.45 This function takes a string as an argument and returns a new string with each character’s case inverted.6 Uppercase letters are converted to lowercase, and lowercase letters are converted to uppercase.78 Solution Plan:9 1. Create a result variable to hold the updated string.10 2. Iterate through each character in the string.11 3. Check if the character is uppercase; if so, convert it to lowercase and add it to the result.12 4. If the character is lowercase, convert it to uppercase and add it to the result.13 5. After iterating through all characters, return the result.14 """ -
Detail Expansion: Besides the original question, provide a few key important details or explanations without which is hard to solve the question.
Detail Elaboration: Provide original question along with the toolbox (commonsense knowledge) to solve the question.
@@
Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes. Each box has outside dimensions of 5 inches by 6 inches by 4 inches. The walls of each box are 1 inch thick, uniformly throughout each face of the box, thereby reducing the inner dimensions of each box. The material of the boxes is uniformly distributed and does not bulge or cave in thereby affecting the inner volume. There are no internal structures or partitions inside the boxes that could further reduce the inner volume. What is the total inner volume of all 3 boxes? Example Detail: Besides providing the input and output of each example, it also provide a step by step explanation of how the input is transformed to the output. Changed from G.3: Generate a python function that fulfills the requirement in docstring and examples usages below.1def derivative(xs: list):2 """ xs represent coefficients of a polynomial.3 xs[0] + xs[1] * x + xs[2] * x^2 + ....4 Return derivative of this polynomial in the same form.56 >>> derivative([3, 1, 2, 4, 5]) calculates the derivative as [1*1, 2*2, 3*4, 4*5] resulting in [1, 4, 12, 20].78 >>> derivative([1, 2, 3]) calculates the derivative as [1*2, 2*3] resulting in [2, 6].9 """
@@
Restrict Question: Adding a new piece of information that serves as a constraint or modifier on the query. Example:
@@
Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes, each of which is externally measured as 5 inches by 6 inches by 4 inches. The boxes have walls that are 1 inch thick. There is also an added wooden board divider in the middle across the smallest dimension which is also 1 inch thick. What is the total inner volume of all the boxes? Restrict Requirement: Add a piece of information that serves as a constraint or modifier on the function.
Changed from G.11def flip_case(string: str, index: int) -> str:23 """For a given string, flip lowercase characters to uppercase and uppercase to lowercase. Only flip the case for characters at indices which are multiples of the provided index.4 Note: If the index provided is 2, only the characters at the 2nd, 4th, 6th positions and so on will have their cases flipped.56 >>> flip_case(’Hello’, 2)7 ’HeLlO’8 """ -
G9.
Subsequent Question: Adding an additional query or requirement based on the answer of of the original question.
@@
Further Question: Adding an additional query that will need extra steps of calculation based on the final answer of the original query. {mdframed}[backgroundcolor=pink!20] Changed from F.2: John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. What is the total inner volume of all 3 boxes? If John wants to entirely fill these boxes with small cubes each measuring 0.5 inches on all sides, then how many cubes will he need? Further Requirement: Adding an additional requirement of transformation based on the output of the original function.1def flip_case_count(string: str) -> Tuple[str, int]:23 """4 For a given string, flip lowercase characters to uppercase and uppercase to lowercase. Additionally, return the number of case flips performed.56 >>> flip_case_count(’Hello’)7 (’hELLO’, 5)8 """ -
G10.
Concurrent Question: Adding an additional query or requirement that is independent from the original question.
@@
Parallel Question: Adding an additional query along with the original query based on the information given in the question, the added query should inquiry a value that is irrelevant of the original answer. Example:
Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. What is the total inner volume of all 3 boxes? What is the total volume of the material used to build the boxes?
Parallel Requirement: Adding an additional requirement in Docstring that does not rely on the output of the original question.
Changed from G.1:1def flip_case_and_count(string: str) -> Tuple[str, int]:23 """For a given string, not only should you flip lowercase characters to uppercase and uppercase to lowercase. You should also output another Title case where only the first letter of each word is capitalized"""45 """>>> flip_case_and_count(’Hello’)6 (’hELLO’, ’Hello’)7 """ -
G11.
Change Question: Change the current query or requirement to a different but similar one based on the existing information provided inside the question.
Change Query: Change the query to ask for another value that requires more computation based on the information given in the question.
@@
Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. What is the total outer volume of all 3 boxes? Change Docstring: Change the Docstring to another requirement based on the input given in the question header.
Changed from G.3:1def calc_derivative(xs: list):23 """ xs represent coefficients of a polynomial.4 xs[0] * (exp (x))^0 + xs[1] * (exp(x))^1 + xs[2] * (exp(x))^2 + ....5 Return derivative of this polynomial in the same form.6 >>> derivative([3, 1, 2, 4, 5])7 [1, 4, 12, 20]8 >>> derivative([1, 2, 3])9 [2, 6]10 """ -
G12.
Info Recombination: Combine the fundamental concepts or frameworks from another question with the original question.
Info Recombination (Math): Graft mathematical structure from another question and combine with the original question.
@@
Changed from F.4: {mdframed}[backgroundcolor=pink!20] Vicki and James are planning an event at their high school that combines a pop singing concert and dance events. The whole event will be 2 hours long. Vicki is allowing each musical group 2 minutes to get on stage, 6 minutes to perform, and then 2 minutes to exit the stage. James will also perform two solo dance routines, each lasting five minutes. Considering a 10-minute intermission during the show, how many musical groups can perform at the concert? Info Recombination (Code): Merge the requirement from another coding question with existing question. Changed from G.1:1def flip_case_and_odd_sum(string: str) -> tuple:2 """3 Given a string, flip lowercase characters to uppercase and uppercase to lowercase.4 Also return the odd letters that are in even positions of the original string.5 string Index starts from 0, alphabet index start from 1. Aa is 1, Bb is 2..6 Examples:7 >>> flip_case_and_odd_sum(’Hello’)8 (’hELLO’, ’o’)9 """ -
G13.
Domain Knowledge: Introduce a specific knowledge in math or code and merge it with the question.
Theoretical Challenge: Incorporate a specific theorem into the question so that perturbed question requires a new toolbox to solve.
@@
Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has an infinite number of boxes numbered as first, second, third, and so on. The first box is 5 inches by 6 inches by 7 inches in size. Starting from the second box each box is half the size of the previous box in each dimension. What is the total volume all the boxes combined? Code Import: The requirement requires to use a specific python library to solve the problem. Changed from G.1: Rewrite the function below to take in batch input parameters and use the multicore cpu for efficiency. -
G14.
Complex Reality: Add an aspect of complexity in the real world scenario.
Value Probability:Introduce concept of uncertainty to deterministic values and calculate the estimation. The perturbed question will require toolbox (knowledge) of probability.
@@
Changed from F.1: {mdframed}[backgroundcolor=pink!20] A merchant wants to make a choice of purchase between 2 purchase plans: jewelry worth $5,000 or electronic gadgets worth $8,000. His financial advisor speculates that the jewelry market has a 70% chance to go up 2.5% and a 30% chance to remain the same, while the electronic gadgets market will rise 1.2% within the same month. If the merchant is looking to maximize profit at the end of this month by making a choice, how much estimated profit would this be? Example Boundary: Add boundary examples along with the existing examples. The boundary examples contains input that does not met requirement specified in the docstring. Changed from G.3: Write a function to fulfill the requirement and all the examples inside the docstring1def derivative(xs: list):23 """ xs represent coefficients of a polynomial.4 xs[0] + xs[1] * x + xs[2] * x^2 + ....5 Return derivative of this polynomial in the same form. The solution should pass all the test cases specified below67 # Regular case8 >>> derivative([3, 1, 2, 4, 5])9 [1, 4, 12, 20]10 # Smaller case11 >>> derivative([1, 2, 3])12 [2, 6]13 # Special case with empty list14 >>> derivative([])15 []16 # Boundary case, the shortest polynomial17 >>> derivative([1])18 [0]19 # Boundary case, all-zero polynomial20 >>> derivative([0.0, 0.0, 0.0])21 [0, 0]22 """ -
G15.
General Solution: Provide the solution in a more general scenario.
Code Implementation: Develop a code function to solve the question in general.
@@
Changed from F.2: {mdframed}[backgroundcolor=pink!20] # Original Examples # Can you write a Python code to find out what is the total inner volume of all 3 boxes?
Higher Order: Write a higher order function that can solve the coding problem in general.
Changed from G.21def greatest_common_divisor(numbers: list[int]) -> int:2 """3 Calculates the greatest common divisor (GCD) of a list of integers.4 Returns the GCD as an integer.56 Examples:7 - For numbers = [20, 40, 60], the function returns 20.8 - For numbers = [35, 14], the function returns 7.9 """
@@
Value Big: Significantly increasing the magnitude of values that pose a challenge for calculations.
Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3000 boxes. Each box is 500 inches by 600 inches by 400 inches. The walls are 100 inches thick. What is the total inner volume of all the boxes? Generalize Parameter: Extend the current parameter into different python object types
Changed from G.2:1def find_common_divisor(value1: Union[int, float, str], value2: Union[int, float, str]) -> float:2 """3 Takes two values (int, float, or float in string format) and finds the largest float that divides both into integers.4 Inputs can be a mix of types. Returns the divisor as a float.56 Examples:7 print(find_common_divisor("0.5", 1)) # 0.58 print(find_common_divisor(0.25, "1.25")) # 0.259 """ -
G17.
Change Value: Change the content of the value to a different one.
Change Subject: If there are multiple mentions in the question, Exchange values of names or references in the question.
Changed from F.5: {mdframed}[backgroundcolor=pink!20] Together David, Bodhi, and Lily collected 43 insects. David found 7 more than Bodhi. Bodhi found half of what Lily found. How many insects did Lily find? Parameter Content: Change the format or meaning of the input parameter.
Changed from G.3:1def derivative(polynomial: str):23 """ ’polynomial’ is a string that stands for polynomial for form4 coefficients_0 + coefficients_1 * x + coefficients_2 * x^2 + ....5 This function will return the derivative of the aforementioned polynomial in the same format.67 >>> derivative(’3 +1x + 2x^2 + 4x^3 + 5x^4’)8 ’1 + 4x + 12x^2 + 20x^3’9 >>> derivative(’1 - 2x + 3x^2’)10 ’-2 + 6x’11 """ -
G18.
Change Operation: Change one operation regarding how the Values are processed.
Change Calculation: Change no more than 3 words in original question so that the toolbox (mathematical operations) involved in the calculation are changed.
Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes. The inner dimension of each box is 3 inches by 4 inches by 2 inches. The walls are 0.5 inches thick. What is the total outer volume of all 3 boxes? Variable Type: Change the python object type of the original parameter while keep its content the same, also specify the return variable to be in a specific type.
Changed from G.3:1def derivative(xs: list[str]) -> list[str]:23 """ xs represent coefficients of a polynomial.4 xs[0] + xs[1] * x + xs[2] * x^2 + ....5 Return derivative of this polynomial in the same form.6 """
Formulation Adjustment: This dimension tests the abstract reasoning ability of LLMs under the same logical structure of the original question. This dimension focuses on solving the general version of the original reasoning problem, rather than focusing on to get a standard solution. For math, We change the context to include one or more symbolic variables to replace its original values.
@@
Variable Response: Replace one value inside the question with a variable and answer with the variable included.